# Cosmic superstrings, metastable strings and ultralight primordial black holes: from NANOGrav to LIGO and beyond

### v1

@@ Abstract
While topologically stable cosmic strings are disfavored by the recent observation of nHz stochastic gravitational waves (GW) by Pulsar Timing Arrays (PTA), e.g., NANOGrav, cosmic metastable strings and superstrings are not. However, because the gravitational waves from all classes of strings generally span a wide range of frequencies, they contradict LIGO’s non-observation of stochastic gravitational waves at the $f\sim$ 25 Hz band for a substantial string-parameter space favoured by the PTA data. Suppose ultralight primordial black holes ($M_{BH}<10^{9}$ g) existed in the early universe. In this case, they reduce the amplitude of the GWs at higher frequencies by providing an early matter-dominated phase, alleviating the tension between LIGO observation and PTA data. We show that the recent PTA data complemented by future LIGO-Virgo-KAGRA (LVK) runs plus detectors such as LISA and ET would be able to dapple the properties and further search strategies of such ultralight primordial black holes which are otherwise fairly elusive as they evaporate in the early universe by Hawking radiation.
@@ I Introduction
Ultralight primordial black holes PBHs [1, 2, 3] with initial mass $M_{\rm BH}\lesssim 10^{8}\,g$ are elusive objects as they evaporate in the early universe, therefore, they can not be constrained by astrophysical means like heavier black holes [4, 5]. Nonetheless, if they existed abundantly and dominated the energy density in the early universe, plenty of direct or indirect observable signatures are achievable in the form of gravitational radiation. For instance, they emit gravitons to constitute a high-frequency gravitational wave background [6, 7, 8, 9] and produce density fluctuation to induce GW background testable in the planned GW detectors [10, 11, 12, 13, 14]. In addition, a PBH domination imprints the spectrum of stochastic gravitational waves background (SGWB) originating from other independent sources. Such imprints on the GWs could be probed with the spectral shape reconstruction techniques [15] planned for the GW detectors. Therefore, a combined study of the GW spectrum associated with the ultralight PBHs–the GW spectrum obtained directly from PBHs plus inquiring into the imprints of a PBH-dominated phase on an independent GW spectrum could be a riveting way to understand the properties of ultralight PBHs. The latter case is the central theme of this article. We study imprints of PBHs on GW background produced by cosmic superstrings [16, 17, 18, 19, 20, 21, 22, 23, 24] and metastable strings [25, 26, 27, 28, 29, 30, 31, 32, 33] [^1] Effect of PBHs on GWs from stable cosmic strings have been recently explored in Refs.[34, 35]. We classify the other GW backgrounds obtained directly from PBH evaporation or their dynamics in the early universe as complementary probes that could be envisaged to distinguish the effect of PBH domination on GWs from cosmic strings from any other early matter domination (EMD) scenarios, see, e.g., [24, 36].
@@
The motivation to consider the mentioned classes of strings is twofold. First, because the ultralight PBHs evaporate at early times, before the Big Bang Nucleosynthesis (BBN) [37, 38, 39], they generically affect any propagating GW spectrum during the PBH-domination phase at higher frequencies. Cosmic strings radiate GWs spanning a wide range of frequencies; generally with a strong scale-invariant amplitude at higher frequencies. Therefore, ultralight PBHs leave possibly the cleanest detectable imprints on such GW spectrum. In addition, a wide-spanning GW signal has the following advantage: if a measurement at a certain frequency band fixes the parameter space describing the GW source, signals at other frequency bands get automatically constrained. This redirects us to the second motivation. Cosmic superstrings and metastable strings are speculated to be among the possible sources of nHz SGWB recently discovered by the Pulsar Timing Arrays (PTA)  [40, 41, 42, 43, 44, 45]. Unlike stable cosmic strings [46, 47, 48, 49] described only by the parameter: $G\mu$, where $G$ is the Newton’s constant and $\mu\sim v_{\Phi}^{2}$ is the string tension with $v_{\Phi}$ being the vacuum expectation value of the scalar field describing the strings  [50, 51, 52], cosmic superstrings and metastable cosmic strings fit the data better [24, 53, 44, 54, 55, 56, 57, 58, 59, 60] because the last two being a two-parameter string model. In addition to $G\mu$, the inter-commutation probability $P$ for superstrings and the parameter $\kappa$ describing the separation between the Grand Unified Theory (GUT) scale and the string scale for metastable strings provide additional degrees of freedom to match the required GW amplitude and spectral slope at nHz frequencies. On the other hand, the high-frequency scale-invariant behaviour of the GW spectrum remains similar for all string classes. Fitting PTA data with GWs from superstrings and metastable strings predicts extremely strong scale-invariant amplitude at higher frequencies if the universe undergoes standard cosmological expansion. In fact, the high-frequency amplitudes are so strong that a significant string parameter space consistent with the PTA data gets ruled out by the LIGO-O3 run which did not observe any SGWB [61]. This puts superstring and metastable string interpretation to the PTA data perhaps in jeopardy unless an additional GW contribution from supermassive black holes (SMBH) is considered to relax the tension somewhat [44]. A PBH-dominated phase in the early universe, however, can cause a characteristic fall-off of the spectrum at higher frequencies, alleviating the tension with the PTA and LIGO completely. Besides, because the PTA data constrains the string parameters and the PBH mass is bounded from above so that they evaporate before BBN, even though reduced, the predicted overall signal strength could be pretty strong.
@@
Focusing exclusively on the parameter range of superstrings and metastable strings favoured by the PTA data, and considering non-rotating PBHs with monochromatic initial mass, we find that PBHs with an initial mass within the range $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$ would imprint the GWs at higher frequencies so that they evade the LIGO bound. We broadly classify such viable signals into two classes: whether or not the next LIGO run (LVK-Design) [62] could probe the PBH-imprinted GW signal obtained from superstrings and metastable strings. We find that for most of the mentioned mass range, PBHs can produce signals at LVK-D while being consistent with the PTA data. Heavier PBHs with an initial mass around $10^{8}\,g$ reduce the GW amplitude such that the signals fall short of the sensitivity reach of LVK-D, however strong enough to produce a Signal-to-Noise-Ratio at ET [63] as large as $\sim 10^{5}$.
@@
In the companion theory paper by NANOGrav collaboration [44], the effect of an EMD on GWs from superstrings and metastable strings has been mentioned to evade the LIGO-O3 bound. Therefore, root-idea-wise, this article does not offer anything new. We rather report if the PBHs provide such EMD, the scenario becomes pretty distinct compared to any other EMD, e.g., provided by a scalar field. This is because, first, the PBHs provide additional GW signatures. For the obtained mass range $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$, although the GW in the form of gravitons appears at extremely high frequency, GWs, e.g., induced by PBH density fluctuation appear in the detectable range. It is not trivial to reconstruct such a combined GW signal by any other means. Second, lately, there has been an extensive effort to understand the effects of ultralight PBHs on Beyond the Standard Model (BSM) physics such as particle dark matter (DM) production and baryon asymmetry of the universe (BAU), see, e.g., Refs.[64, 65, 66, 67] (subsequent references are in Sec.IV.2). For the $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$, such effects could be quite significant, specifically, BSM physics occurring around the Electroweak (EW) scale, even down to the BBN (MeV) scale. In this context, we discuss a recently proposed BSM model [68, 69] accommodating neutrino mass, BAU via leptogenesis, super-heavy DM, and metastable strings (the last feature is new). We show how a PBH-string landscape could be capable of probing untrod parameter regions of BSM models.
@@ II Stable strings, superstrings, metastable strings, and non-standard cosmology
In Eq.(II.3), $\beta=2/3(1+\omega)$ with $\omega$ is the equation of state parameter of the universe, and $A_{\beta}=5.4$ ($A_{\beta}=0.39$) for radiation-dominated (matter-dominated) universe [80]. The quantity $\Gamma_{k}={\Gamma k^{-\delta}}/{\zeta(\delta)}$ quantifies the emitted power in $k$-th mode, with $\delta=4/3$ ($\delta=5/3$) for loops containing small-scale structures such as cusps (kinks) [81] and makes it evident that $k=1$ (fundamental) mode is the dominant one contributing to the GWs. In principle, the GW spectrum corresponding to the fundamental mode broadly captures most of the underlying features of a given BSM model producing cosmic strings. However, when it comes to comparing the spectrum with real data, e.g., PTA data, and based on that, one intended to forecast future predictions, it is desirable to sum a large number of modes to obtain accurate results.
@@
Without any intermediate matter-dominated epoch, Eq.(II.2) can be solved to obtain the GW spectra taking into account two main contributions: i) from the loops created in the radiation domination but emitting in the standard matter domination, ii) from the loops created as well as emitting in the radiation domination. The former corresponds to a low-frequency peak while the latter produces a scale-invariant plateau at high frequencies (see Fig.1) which for $k=1$ mode is given by [75, 76, 85]
@@
where $T_{\rm brk}\equiv T_{\rm end}$ is the temperature corresponding to the end of early matter domination. When summed over a large number of modes ($k\rightarrow\infty$), two important changes occur in the GW spectrum. First, the plateau amplitude increases approximately by a factor of $\zeta(\delta)$, i.e., $\Omega_{\rm GW}^{\rm plt,\infty}\simeq\Omega_{\rm GW}^{\rm plt}\zeta(\delta)$ and the spectrum falls as $f^{1-\delta}$ instead of $f^{-1}$ [87, 49].
@@
Figure 1: All GW spectra are produced for $k=1$ mode. Top-left: GW spectra for cosmic super (stable) strings shown with green solid (dashed) curve. Two benchmark string parameters chosen for superstrings are $P=6\times 10^{-3}$ and $G\mu=2\times 10^{-12}$. For the stable strings, we rescale $G\mu$ as $G\mu\rightarrow G\mu/P^{2}$ so that comparable plateau amplitude is achieved in both cases. The horizontal and vertical dotted black lines represent the analytical expression for the plateau amplitude (multiplied by $P^{-1}$ for superstrings), and the spectral break frequency $f_{\rm brk}$ corresponding to a long-lasting matter domination ending at $T=0.1$ GeV. The green vertical lines correspond to the spectral break frequencies where the spectrum would have fallen due to particle production if there were no matter domination. Top-right: GW spectra for cosmic metastable (stable) strings are shown with the red solid (dashed) curve. In this case, two benchmark string parameters are chosen as $\sqrt{\kappa}=8$ and $G\mu=2\times 10^{-7}$. The dotted horizontal and vertical lines represent the same as in the left panel figure. The red vertical line represents the frequency at which the spectrum would have fallen if there was no matter domination. The GW spectrum represented by the blue curve is for $\sqrt{\kappa}=6$ assuming no matter domination. The density reach of SKA [92], LISA [93], DECIGO [94] and LVK-D [62] are shown with the filled coloured curves. Bottom: The benchmark GW spectra in the PTA band with the recent posteriors reported by NANOGrav recently.
@@
We conclude the discussion by showing the GW spectra in Fig.1 for superstrings and metastable strings comparing them to the ones from stable strings. We choose a benchmark value of $T_{\rm brk}=0.1$ GeV and consider only the $k=1$ mode, leaving a more realistic analysis by summing a large number of modes for the next section. In Fig.1 (left) we show the GW spectrum for cosmic super (stable) strings with the green solid (dashed) curve. For the superstrings, we choose the benchmark values $P=6\times 10^{-3}$ and $G\mu=2\times 10^{-12}$ while for the stable cosmic strings, we rescale $G\mu$ as $G\mu\rightarrow G\mu/P^{2}$ such that comparable plateau amplitude can be obtained in both cases. Note that, for superstrings, one has the required blue-tilted GW spectrum in the PTA band (A blue-tilted distribution of fourteen NANOGrav posteriors is shown in black; in this work, we use only NANOGrav data which are consistent with all other PTAs (IPTA) [95] ). On the other hand, even though the required GW amplitude can be obtained for the stable strings, the spectrum is nearly scale-invariant in the PTA band and thereby does not fit the data. A similar situation occurs in the case of metastable strings as shown in the top right panel with the red curves [^2] With the blue curve we show the GW spectra for $\sqrt{\kappa}=6$ evading all the bounds with the dominant contribution at very high-frequencies. Although this spectrum is not so relevant for the present study, we find it interesting to mention it as it is generally not present in the case of stable (and super) strings.. In both plots, the black dotted lines represent the analytical expression for the plateau amplitude and the spectral break frequency given in Eq.(II.4) and (II.5), respectively. The green and the red vertical lines represent the spectral break frequency $f_{\rm cusp}$ at which the spectrum would have fallen if there was no matter domination. In the bottom panel, we zoomed in on the GW spectra within the PTA band to show the consistency of the blue-tiled GW spectrum obtained from cosmic superstrings and metastable strings with the data.

### v2

@@ Abstract
While topologically stable cosmic strings are disfavoured by the recent observation of nHz stochastic gravitational waves (GW) by Pulsar Timing Arrays (PTA), e.g., NANOGrav, cosmic metastable strings and superstrings are not. However, because the gravitational waves from all classes of strings generally span a wide range of frequencies, they contradict LIGO’s non-observation of stochastic gravitational waves at the $f\sim$ 25 Hz band for a substantial string-parameter space favoured by the PTA data. Suppose ultralight primordial black holes ($M_{\rm BH}<10^{9}$ g) existed in the early universe. In this case, they reduce the amplitude of the GWs at higher frequencies by providing an early matter-dominated phase, alleviating the tension between LIGO observation and PTA data. We show that the recent PTA data complemented by future LIGO-Virgo-KAGRA (LVK) runs plus detectors such as LISA and ET would be able to dapple the properties and further search strategies of such ultralight primordial black holes which are otherwise fairly elusive as they evaporate in the early universe by Hawking radiation.
@@ I Introduction
Ultralight primordial black holes PBHs [1, 2, 3] with initial mass $M_{\rm BH}\lesssim 10^{8}\,g$ are elusive objects as they evaporate in the early universe, therefore, they can not be constrained by astrophysical means like heavier black holes [4, 5]. Nonetheless, if they existed abundantly and dominated the energy density in the early universe, plenty of direct or indirect observable signatures are achievable in the form of gravitational radiation. For instance, they emit gravitons to constitute a high-frequency gravitational wave background [6, 7, 8, 9] and produce density fluctuations to induce GW background testable in the planned GW detectors [10, 11, 12, 13, 14]. In addition, a PBH domination imprints the spectrum of stochastic gravitational waves background (SGWB) originating from other independent sources. Such imprints on the GWs could be probed with the spectral shape reconstruction techniques [15] planned for the GW detectors. Therefore, a combined study of the GW spectrum associated with the ultralight PBHs–the GW spectrum obtained directly from PBHs plus inquiring into the imprints of a PBH-dominated phase on an independent GW spectrum could be a riveting way to understand the properties of ultralight PBHs. The latter case is the central theme of this article. We study imprints of PBHs on GW background produced by cosmic superstrings [16, 17, 18, 19, 20, 21, 22, 23, 24] and metastable strings [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] [^1] Effect of PBHs on GWs from stable cosmic strings have been recently explored in Refs.[36, 37]. We classify the other GW backgrounds obtained directly from PBH evaporation or their dynamics in the early universe as complementary probes that could be envisaged to distinguish the effect of PBH domination on GWs from cosmic strings from any other early matter domination (EMD) scenarios, see, e.g., [24, 38].
@@
The motivation to consider the mentioned classes of strings is twofold. First, because the ultralight PBHs evaporate at early times, before the Big Bang Nucleosynthesis (BBN) [39, 40, 41, 42], they generically affect any propagating GW spectrum during the PBH-domination phase at higher frequencies. Cosmic strings radiate GWs spanning a wide range of frequencies; generally with a strong scale-invariant amplitude at higher frequencies. Therefore, ultralight PBHs leave possibly the cleanest detectable imprints on such GW spectrum. In addition, a wide-spanning GW signal has the following advantage: if a measurement at a certain frequency band fixes the parameter space describing the GW source, signals at other frequency bands get automatically constrained. This redirects us to the second motivation. Cosmic superstrings and metastable strings are speculated to be among the possible sources of nHz SGWB recently discovered by the Pulsar Timing Arrays (PTA)  [43, 44, 45, 46, 47, 48]. Unlike stable cosmic strings [49, 50, 51, 52] described only by the parameter: $G\mu$, where $G$ is the Newton’s constant and $\mu\sim v_{\Phi}^{2}$ is the string tension with $v_{\Phi}$ being the vacuum expectation value of the scalar field describing the strings  [53, 54, 55], cosmic superstrings and metastable cosmic strings fit the data better [24, 56, 47, 57, 58, 59, 60, 61, 62, 63, 64] because the last two being a two-parameter string model. In addition to $G\mu$, the inter-commutation probability $P$ for superstrings and the parameter $\kappa$ describing the separation between the Grand Unified Theory (GUT) scale and the string scale for metastable strings provide additional degrees of freedom to match the required GW amplitude and spectral slope at nHz frequencies. On the other hand, the high-frequency scale-invariant behaviour of the GW spectrum remains similar for all string classes. Fitting PTA data with GWs from superstrings and metastable strings predicts extremely strong scale-invariant amplitude at higher frequencies if the universe undergoes standard cosmological expansion. In fact, the high-frequency amplitudes are so strong that a significant string parameter space consistent with the PTA data gets ruled out by the LIGO-O3 run which did not observe any SGWB [65]. This puts superstring and metastable string interpretation to the PTA data perhaps in jeopardy unless an additional GW contribution from supermassive black holes (SMBH) is considered to relax the tension somewhat [47]. A PBH-dominated phase in the early universe, however, can cause a characteristic fall-off of the spectrum at higher frequencies, alleviating the tension with the PTA and LIGO completely. Besides, because the PTA data constrains the string parameters and the PBH mass is bounded from above so that they evaporate before BBN, even though reduced, the predicted overall signal strength could be pretty strong.
@@
Focusing exclusively on the parameter range of superstrings and metastable strings favoured by the PTA data, and considering non-rotating PBHs with monochromatic initial mass, we find that PBHs with an initial mass within the range $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$ would imprint the GWs at higher frequencies so that they evade the LIGO bound. We broadly classify such viable signals into two classes: whether or not the next LIGO run (LVK-Design) [66] could probe the PBH-imprinted GW signal obtained from superstrings and metastable strings. We find that for most of the mentioned mass range, PBHs can produce signals at LVK-D while being consistent with the PTA data. Heavier PBHs with an initial mass around $10^{8}\,g$ reduce the GW amplitude such that the signals fall short of the sensitivity reach of LVK-D, however strong enough to produce a Signal-to-Noise-Ratio at ET [67] as large as $\sim 10^{4}$.
@@
In the companion theory paper by NANOGrav collaboration [47], the effect of an EMD on GWs from superstrings and metastable strings has been mentioned to evade the LIGO-O3 bound. Therefore, fundamentally, this article does not offer anything new. We rather report if the PBHs provide such EMD, the scenario becomes very distinct and flexile compared to any other EMD, e.g., provided by a scalar field. This is because, first, as mentioned, the PBHs provide additional GW signatures. For the obtained mass range $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$, although the GW in the form of gravitons appears at extremely high frequency, GWs, e.g., induced by PBH density fluctuations appear in the detectable range. It is not trivial to reconstruct such a combined GW signal by any other means. Second, lately, there has been an extensive effort to understand the effects of ultralight PBHs on beyond the Standard Model (BSM) physics such as particle dark matter (DM) production and baryon asymmetry of the universe (BAU), see, e.g., Refs.[68, 69, 70, 71] (subsequent references are in Sec.IV.2). For the $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$, such effects could be quite significant, specifically, BSM physics occurring around the Electroweak (EW) scale, even down to the BBN (MeV) scale. In this context, we discuss a recently proposed BSM model [72, 73] accommodating neutrino mass, BAU via leptogenesis, super-heavy DM, and metastable strings (the last feature is new). We show how a PBH-string landscape could be capable of probing untrod parameter regions of BSM models.
@@ II Stable strings, superstrings, metastable strings, and non-standard cosmology
In Eq.(II.3), $\beta=2/3(1+\omega)$ with $\omega$ is the equation of state parameter (constant) of the universe, and $A_{\beta}=5.4$ ($A_{\beta}=0.39$) for radiation-dominated (matter-dominated) universe [84]. In this section, for illustrative purposes, we shall treat $A_{\beta}$ as a step function during the change of cosmology plus use Eq.(II.3) to compute gravitational waves spectra. We shall account for the actual time evolution of $A_{\beta}$ with VOS equations [81, 82, 83, 84] and appropriate loop number density [47] in the next section. The quantity $\Gamma_{k}={\Gamma k^{-\delta}}/{\zeta(\delta)}$ quantifies the emitted power in $k$-th mode, with $\delta=4/3$ ($\delta=5/3$) for loops containing small-scale structures such as cusps (kinks) [85] and makes it evident that $k=1$ (fundamental) mode is the dominant one contributing to the GWs. In principle, the GW spectrum corresponding to the fundamental mode broadly captures most of the underlying features of a given BSM model producing cosmic strings. However, when it comes to comparing the spectrum with real data, e.g., PTA data, and based on that, one intended to forecast future predictions, it is desirable to sum a large number of modes to obtain accurate results.
@@
Without any intermediate matter-dominated epoch, Eq.(II.2) can be solved to obtain the GW spectra taking into account two main contributions: i) from the loops created in the radiation domination but emitting in the standard matter domination, ii) from the loops created as well as emitting in the radiation domination. The former corresponds to a low-frequency peak, while the latter produces a scale-invariant plateau at high frequencies (see Fig.1) which for $k=1$ mode is given by [79, 80, 89]
@@
where $T_{\rm brk}\equiv T_{\rm end}$ is the temperature corresponding to the end of early matter domination. When summed over a large number of modes ($k\rightarrow\infty$), two important changes occur in the GW spectrum. First, the plateau amplitude increases approximately by a factor of $\zeta(\delta)$, i.e., $\Omega_{\rm GW}^{\rm plt,\infty}\simeq\Omega_{\rm GW}^{\rm plt}\zeta(\delta)$ and the spectrum falls as $f^{1-\delta}$ instead of $f^{-1}$ [91, 92, 52, 93].
@@
Figure 1: All GW spectra are produced for $k=1$ mode. Top-left: GW spectra for cosmic super (stable) strings shown with green solid (dashed) curve. Two benchmark string parameters chosen for superstrings are $P=6\times 10^{-3}$ and $G\mu=2\times 10^{-12}$. For the stable strings, we rescale $G\mu$ as $G\mu\rightarrow G\mu/P^{2}$ so that comparable plateau amplitude is achieved in both cases. The horizontal and vertical dotted black lines represent the analytical expression for the plateau amplitude (multiplied by $P^{-1}$ for superstrings), and the spectral break frequency $f_{\rm brk}$ corresponding to a long-lasting matter domination ending at $T=0.1$ GeV. The green vertical lines correspond to the spectral break frequencies where the spectrum would have fallen due to particle production if there were no matter domination. Top-right: GW spectra for cosmic metastable (stable) strings are shown with the red solid (dashed) curve. In this case, two benchmark string parameters are chosen as $\sqrt{\kappa}=8$ and $G\mu=2\times 10^{-7}$. The dotted horizontal and vertical lines represent the same as in the left panel figure. The red vertical line represents the frequency at which the spectrum would have fallen if there was no matter domination. The GW spectrum represented by the blue curve is for $\sqrt{\kappa}=6$ assuming no matter domination. The density reach of SKA [98], LISA [99], DECIGO [100] and LVK-D [66] are shown with the filled coloured curves. Bottom: The benchmark GW spectra in the PTA band with the recent posteriors reported by NANOGrav.
@@
We conclude the discussion by showing the GW spectra in Fig.1 for superstrings and metastable strings comparing them to the ones from stable strings. We choose a benchmark value of $T_{\rm brk}=0.1$ GeV and consider only the $k=1$ mode, leaving a more realistic analysis by summing a large number of modes for the next section. In Fig.1 (left) we show the GW spectrum for cosmic super (stable) strings with the green solid (dashed) curve. For the superstrings, we choose the benchmark values $P=6\times 10^{-3}$ and $G\mu=2\times 10^{-12}$ while for the stable cosmic strings, we rescale $G\mu$ as $G\mu\rightarrow G\mu/P^{2}$ such that comparable plateau amplitude can be obtained in both cases. Note that, for superstrings, one has the required blue-tilted GW spectrum in the PTA band (A blue-tilted distribution of fourteen NANOGrav posteriors is shown in gray; in this work, we use only NANOGrav data which are consistent with all other PTAs (IPTA) [101]). On the other hand, even though the required GW amplitude can be obtained for the stable strings, the spectrum is nearly scale-invariant in the PTA band and thereby does not fit the data. A similar situation occurs in the case of metastable strings as shown in the top right panel with the red curves [^2] With the blue curve we show the GW spectra for $\sqrt{\kappa}=6$ evading all the bounds with the dominant contribution at very high-frequencies. Although this spectrum is not so relevant for the present study, we find it interesting to mention it as it is generally not present in the case of stable (and super) strings.. In both plots, the black dotted lines represent the analytical expression for the plateau amplitude and the spectral break frequency given in Eq.(II.4) and (II.5), respectively. The green and the red vertical lines represent the spectral break frequency $f_{\rm cusp}$ at which the spectrum would have fallen if there was no matter domination. In the bottom panel, we zoomed in on the GW spectra within the PTA band to show the consistency of the blue-tiled GW spectrum obtained from cosmic superstrings and metastable strings with the data.